A 400. mL sample of a gas at 10.0°C is warmed to 25.0°C at a constant pressure. Calculate the final volume
1
Expert's answer
2020-06-04T10:38:52-0400

According to Charles's law, or the law of volumes, the ratio of the volume to the temperature is constant at constant pressure and mass of an ideal gas:

"\frac{V}{T} = const" .

Therefore, if the initial volume is 400 mL and the temperature changes from 10°C to 25°C, the final volume is:

"\frac{V_1}{T_1} = \frac{V_2}{T_2}; V_2 = \frac{V_1T_2}{T_1} = \frac{400·(25.0+273.15)}{10.0 + 273.15} = 421" mL.

Note: the volume is proportional to the absolute temperature. Don't forget to convert from celsius to kelvin ( add 273.15).

Answer: the final volume is 421 mL.